find the equation of a line whose y- intercept is (0,0) and whose slope is 4?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

find the equation of a line whose y- intercept is (0,0) and whose slope is 4?
-----
y = mx+b
m = slope = 4
b = intercept = 0
---
Equation:
y = 4x
=============
Cheers,
Stan H.
=======

Answer by mananth(16946)

m= 4

Plug value of the slope and point ( 0 , 0 ) in
Y = m x + b
0.00 = 0 + b
b= 0 - 0
b= 0
So the equation will be
Y = 4 x
m.ananth@hotmail.ca